If the energy levels in the neon atom are 4.3 x 10 -18 J and 3.97 x 10 -18 J and 1.07 x 10 -18 J Use the data to calculate two photon frequencies that could be absorbed by neon in the lowest state
1
Expert's answer
2019-08-19T09:18:19-0400

photon energy is the difference in level energies , which means the frequency "\nu = \frac{\Delta E}{h}"


consequently "\nu _{1} = 10^{16} \cdot\frac{4.3-1.07}{6.63} = 0.
49\cdot10^{16} Hz"


and "\nu _{2} = 10^{16} \cdot\frac{3.97-1.07}{6.63} = 0.44\cdot10^{16} Hz"